# LiDDA: Data Driven Attribution at LinkedIn

### v1

@@ 2.5. Imputation of Paid Media Touchpoints
When a member clicks on an ad on an external platform they land on a LinkedIn microsite. We can use the URLs from microsites as a proxy for paid clicks at the member level. On the other hand, we do not have any impression data at the member level. The lack of member-level impression data for paid media channels can bias attribution modeling.
While member-level data is not available for paid media channels, we can still obtain daily aggregate campaign-level data, e.g, number of impressions for a campaign on a given day, from external platforms. We leverage this aggregated data to mitigate the bias by imputing paid media impressions probabilistically. This is performed in two steps. First, we attach a certain number of impressions to existing click events, because each click necessarily must have prior impressions. Second, we look at how impressions are distributed across online members without any reference to clicks for each paid media channel. In both cases, we use owned channels as proxies to get prior information on these distributions, as shown in Figure 2.
@@
Algorithm 1 Paid Media Impression Imputation Methodology
Member $touchpoint$ data
Aggregate external (paid) campaign $impression$ data $Imp_{AGG-campaign}$
for $touchpoint$ do
     if $touchpoint$ is a $click$ for $paid-channel$ (ch) then
         Sample $n$ pseudo $impressions$ $Imp_{n}$ from
         $N(imp-click-ratio_{ch},\sigma^{2})$
     end if
end for
Compute total assigned pseudo impressions for each campaign:
   $Imp_{campaign}$ = $\sum_{paid-campaigns}Imp_{n}$
if $I_{AGG-remainder}$ = $\sum(I_{AGG-campaign})$ $Imp_{campaign})$ 0 then
     Sample $N$ random members
     for $member$ do
         Sample $I_{R}\sim\mathcal{N}(\mu,\sigma^{2})$ random impressions
         Create touchpoint $I_{R}$ until $\sum(I_{R})=I_{AGG-remainder}$
     end for
end if
One of the main challenges in leveraging multiple third-party platforms is that they report aggregate impressions for both members and non-members. On the other hand, DDA models only member data. Each 3rd party channel can have a different percentage of identifiable impressions. To this end, we estimate a downscaling factor for different 3rd party platforms using internal click-based based data to determine how many impressions should be distributed among members while modeling.
We analyze the distribution of how many impressions proceed each click event at the member level from owned channels. We also calculate the ratio of daily aggregate impressions to click from owned channels. Using these ratios, daily aggregate impression-to-clicks and the impressions-per-click at the member level, we infer how many impressions should be assigned to each click event for different paid-channels.
Using this prior information and daily impression-to-click ratio for each external paid channel, we assign impressions to existing paid proxy click events sampled from the distribution $N(imp-click-ratio_{ch},\sigma^{2})$. Next, we distribute the remaining daily impressions from paid media campaigns across members randomly based on campaigns’ and members’ regions. While doing this, we apply an estimated member-to-nonmember ratio realizing that aggregated external impressions may belong to both members and non-members. We make sure the distributed impressions do not exceed the estimated limits. This methodology is summarized in Algorithm 1.
[…]
2.6. Sessionization & Downsampling
Another issue in modeling attribution is that certain channels have significantly more touchpoints compared to other channels in member paths due to signal bias. For example, emails may be sent once a week but in-feed ads could have multiple daily exposures. The paths that go into attribution modeling are of fixed length, therefore touchpoints from owned channels may saturate the paths and reduce the representation of paid (external) touchpoints in paths. To make the representation of channels more balanced across paths, we perform two path adjustments which yield more balanced representations of engagements as shown in Figure 4.

### v2

@@ 2.5. Imputation of Paid Media Touchpoints
When a user clicks on an ad on an external platform they land on a LinkedIn microsite. We can use the URLs from microsites as a proxy for paid clicks at the user level. Meanwhile, we can also obtain daily aggregate campaign-level data, e.g, number of impressions or clicks for a campaign on a given day. Together with member-level clicks and impressions from owned channel, we can impute paid media impressions probabilistically.
Let $i$ index members, $d$ days, $c$ channels, $j\in J_{i,c,d}$ external‐click events, and $k\in\mathcal{P}_{d}$ no‐click paths (i.e., path without clicks from external channel). We perform the imputation in three steps (Algorithm 1): (i) Estimate member impression from external total impressions, yielding total ”budget” that needs allocation; (ii) Compute impression per clicks among the clickers $r_{c,d}$ and then draw impression from corresponding geometric distribution; (iii) Assign weights $u_{k}$ for no-click path $k$ based on campaign and location information, then draw Poisson counts of impressions for no-click path to allocate remaining impressions. Further information, including the step-by-step procedure, implementation notes, the equivalent convex‐optimization formulation, and its solution, is provided in Appendix A.1.
2.6. Sessionization
Another challenge in attribution modeling is that some channels generate far more touchpoints than others, creating a signal imbalance. For example, emails may be sent weekly, whereas in-feed ads can appear multiple times per day. Since our attribution paths have a fixed length, owned-channel touchpoints can dominate and crowd out paid (external) exposures. To address this, we apply two path-adjustment steps, sessionization and downsampling, to produce more balanced engagement sequences (Figure 2).
For sessionization, we group impressions from the same campaign on the same day into a single session‐level touchpoint for specific channels. For downsampling, we apply empirically derived, channel–action–specific drop rates to remove excess touchpoints. In both approaches, we recover the attribution credit to grouped/removed events according to the sampling probability. More details can be found in Appendix A.2
[…]
2.7. Output Calibration
MMM provides channel-level attribution and measures marketing incrementality in the context of macroeconomic factors but cannot trace individual member journeys. DDA captures those journeys at the interaction level but lacks certain signals (e.g. radio, events, TV, or privacy-restricted data). Two independent attribution systems may diverge, particularly if they utilize different interactions, and thus a unified approach is needed(Google, [n. d.]; Experimental, [n. d.]).
We introduce a training-time calibration that extract signals from our MMM model to our DDA model to form a unified macro level view for marketer, while still permitting fine-grained, path-level attributions. Let $a_{j}(\theta)$ be the DDA attribution weight for touchpoint $j$, $A^{\rm DDA}_{c}(\theta)=\sum_{j:\mathrm{chan}(j)=c}a_{j}(\theta)$ the total DDA weight for channel $c$, $A^{\rm MMM}_{c}$ the corresponding channel-level attribution from MMM, $\mathcal{L}_{\rm DDA}(\theta)$ the usual DDA training loss, and $\beta\geq 0$ a tuning parameter. We minimize the combined loss:
$\mathcal{L}_{\rm total}(\theta)=\mathcal{L}_{\rm DDA}(\theta)\;+\;\beta\sum_{c}\bigl{(}A^{\rm DDA}_{c}(\theta)-A^{\rm MMM}_{c}\bigr{)}^{2},$
and provide path level calibration in Algorithm 2.
This calibration “teaches” the DDA model, during training, to respect the macro-level attribution from MMM, eliminating the need for any post-hoc rescaling, while preserving detailed, campaign- and engagement-level attribution within each channel. Details on batch level calibration and path level calibration can be found in Appendix A.3
Algorithm 2 Train DDA Model with Path‐Level MMM Calibration
Inputs:
  • Paths dataset $\{(x_{i},y_{i},s_{i})\}_{i=1}^{N}$, where $x_{i}$ are features, $y_{i}$ labels, and $s_{i}$ the sequence of length $L_{i}$
  • MMM attribution $a_{c}^{\rm MMM}$, channel distribution $\pi_{c}^{\rm MMM}$
  • DDA network $g(\cdot;\theta)$ with parameters $\theta$, that output conversion probability $f$, and attention weight $a$.
  • Hyperparameters: calibration weight $\beta$, learning rate $\eta$, epochs $T$, batch size $B$
Output: Trained parameters $\theta$
for $\mathrm{epoch}=1$ to $T$ do
     Shuffle dataset
     for each batch $\mathcal{B}$ of size $B$ do $\triangleright$ Forward pass
         for each path $i\in\mathcal{B}$ do
              Compute DDA attributions $a_{i}(c;\theta)$ for $c=1,\dots,C$
              Compute path-level channel distribution $\pi^{(i)}_{c}$
              Compute path‐level IPW weights $w^{(i)}_{c}=\frac{\pi^{(i)}_{c}}{\pi_{c}^{\rm MMM}}$
              ”Recover” path-level MMM attribution as target
$\tilde{t}^{(i)}_{c}=w^{(i)}_{c}\,a_{c}^{\rm MMM},\quad t^{(i)}_{c}=\frac{\tilde{t}^{(i)}_{c}}{\sum_{d}\tilde{t}^{(i)}_{d}}$
         end for
         Compute base DDA loss
$\mathcal{L}_{\rm DDA}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\ell\bigl{(}f_{i}(\cdot;\theta),y_{i}\bigr{)}$
         Compute calibration loss (MSE or KL):
$\mathcal{L}_{\rm calib}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\sum_{c=1}^{C}\bigl{(}a_{i}(c;\theta)-t^{(i)}_{c}\bigr{)}^{2}$
         Total loss:
$\mathcal{L}=\mathcal{L}_{\rm DDA}\;+\;\beta\,\mathcal{L}_{\rm calib}$
$\triangleright$ Backward and update
         $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}$
     end for
end for
3. Experimentation Results
The goal of the attribution model is to proportionally quantify the impact of each marketing action towards a conversion. Since there is no ground-truth data on the contribution of each action towards a conversion, we perform the following offline and online validations to evaluate the attribution model.
3.1. Offline Validation
In offline validation, we seek ensure the model’s accuracy and stability as these are intuitively desirable properties of any attribution system. Because our offline process only has access to observational data, we primarily focus on typical ML performance as guardrails rather than actual attribution accuracy, which we will discuss in Section 3.2.
3.1.1. Prediction Stability
We first assess the model’s performance using binary classification metrics, aiming for high predictive accuracy. We hold 10% training data for validation and separate the rest into 10 subsets. We retrain an attribution model on each subset, evaluate it on the holdout dataset, and record its performance. The resulting prediction performances converge ($>0.97$ for ROC-AUC and PR-AUC), confirming that the attribution model has high prediction power and exhibits consistent results.
We also evaluate the stability of predicted conversion rate using the Anderson-Darling statistic (Stephens, 1974; Anderson and Darling, 1952). The result shows that the predicted conversion rate are consistent over time. Further information is provided in Appendix B.1.
3.1.2. Attribution Stability
We perform a similar check on our model’s stability in producing attribution results. For each action, we bootstrap sample that action’s weights with replacement for 100 times, calculate the mean of the samples, and analyze the distribution of the 100 means to evaluate the stability of the attribution weights (Figure 3). We observe that the mean weights of each action converge. While this approach does not have an intrinsic absolute threshold, we use these values as guardrails to monitor for potential issues.
Figure 3. Distribution of bias in the bootstrap means of the action weights, calculated as the difference between each bootstrap mean and the actual sample mean. The bias distributions center around 0.
[…]
3.1.3. Ablation on Model Input
We conducted an ablation study to quantify the impact of each model component by training and evaluating DDA model with particular features removed, including member/company embeddings, campaign embeddings, date embeddings, or all embeddings (sequence only). We compared ROC-AUC against the baseline and measured how the channel-level attribution mix deviated (mean squared deviation on percentage) from the full model. Removing campaign or date features caused small AUC drops (–0.1% and –0.3%), while omitting entity embeddings and using sequence alone reduced AUC more substantially (–1.7% and –5.1%). In terms of attribution stability, date removal produced the largest shift (MSD=49.3), with campaign and entity removals and the sequence-only variant showing smaller but non-negligible deviations (MSD=8.5, 14.4, and 11.8, respectively). Details are in Appendix B.3.
3.2. Online Validation
Offline validation does not give us insight into whether the attention weight based attribution are close to actual member behavior. Thus, we run experiments for online validations to check how close our model’s attributions align with real-world marketing efforts. We can either run member level test for for owned channels or geo-level experiments for external paid channels. We conduct incremental lift tests for email channel. In particular, we randomly split members into two groups. We select certain campaigns ($\mathcal{T}^{h}$) to hold out email sending during the experiment in the control group and conduct business as usual in the treatment group. It is important to note that email sending was sparse across both members and time, limited to approximately one email per week to a small set of selected members within the connection graph. Additionally, AA tests were conducted as a basic validation of the experimental setup, and no statistically significant differences in conversion rates were found between groups.
To validate member-level experiment, we simulate a holdout experiment within the treatment group by creating counterfactual conversion paths, where the holdout campaign touchpoints ($t\in\mathcal{T}^{h}$) are omitted. This approach helps estimate what the conversion path would look like had the member been in the control group. Define two conversion probabilities: $P(Y=1|Z=1,S=S_{1})$ with original path $S_{1}$, and $P_{1}(Y=1|Z=1,S=S_{1}-\mathcal{T}^{h})$ with counterfactual path $S_{1}-\mathcal{T}^{h}$, where $Z=1$ index the treatment group. The corresponding attribution of touchpoints $t\in\mathcal{T}^{h}$ can be defined as:
(3) $a({\mathcal{T}^{h}})=\frac{P_{1}(Y=1|S=S_{1})-P_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})}{P_{1}(Y=1|S=S_{1})}$
where, $P_{1}(\cdot)=P(\cdot|Z=1)$. We compare the distribution of DDA attribution against the measured attribution from experiment.
A similar formulation can be constructed using $P_{0}(\cdot)$. We didn’t go with this route for validation, as the simulation of counterfactual paths ($S_{0}+\mathcal{T}^{h}$) by adding touches requires more assumptions such as sequencing and campaigns which may introduce bias and noise.
3.2.1. Inverse Propensity Weighting
From the experimentation data, we can calculate $\hat{P}_{1}(Y=1|S=S_{1})$ and $\hat{P}_{0}(Y=1|S=S_{1}-\mathcal{T}^{h})$ easily, by simply counting percentage of conversion in the treatment and control group. In order to obtain estimate of $P_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})$, we use inverse propensity weighting (IPW)(Hernán and Robins, 2006). Denote $e_{i}$ as the propensity score for member $i$ and $Z_{i}$ as the indicator for member $i$ receiving campaigns as in the treatment group (=1) and for being in the control group (=0). Denote $E_{M_{i}}$ as the member embeddings for member $i$ that is computed by LinkedIn.
Step 1: Estimate each member’s treatment propensity, and corresponding IPW
(4) $e_{i}=P(Z_{i}=1|E_{M_{i}}),\qquad w_{i}=\begin{cases}1,&\text{if }Z_{i}=1\\ \frac{e_{i}}{1-e_{i}},&\text{if }Z_{i}=0\end{cases}$
Step 2: Compute the attribution
(5) $\hat{a}^{\exp}(\mathcal{T}^{h})=\frac{\hat{P}_{1}(Y=1|S=S_{1})-\hat{P}_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})}{\hat{P}_{1}(Y=1|S=S_{1})}$
by plugging in $\hat{P}_{1}(Y=1|S=S_{1})=(\sum_{i:Z_{i}=1}Y_{i})/(\sum_{i:Z_{i}=1}1)$, and
(6) $\hat{P}_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})=\sum_{i:Z_{i}=0}\frac{w_{i}Y_{i}}{\sum_{i:Z_{i}=0}w_{i}}.$
For validation, we then compute the attention based attribution, i.e.,
(7) $\hat{a}^{DDA}(\mathcal{T}^{h})=\frac{\sum_{i:Z_{i}=1}\sum_{t\in\mathcal{T}^{h}}a_{i}^{DDA}(t)}{\sum_{i:Z_{i}=1}1}$
where $a_{i}^{DDA}(t)$ is attention based attribution for member $i$ for touchpoint $t$. Further details, including validation of IPW and two additional attribution measures are available in Appendix B.4.
3.2.2. Experiment Results
We chose three incrementality tests in the email channel to validate the attribution model and the primary metric is attribution defined in eq. [B.4]. Our goal is to evaluate the effectiveness of the acquisition email series. The workflow is illustrated in Figure 4 and the comparisons are listed in Table 2.
[…]
Test index $\hat{a}^{\exp}$ $\hat{a}^{DDA}$ $\hat{d}^{DDA}$
1 9.68% (1.47%, 20.02%) 7.65% (4.95%, 10.07%) -2.23% (-5.85%, 0.06%)
2 14.05% (3.35%, 24.75%) 14.96% (11.78%, 18.24%) 0.95% (-0.89%, 2.25%)
3 12.54% (2.10%, 27.69%) 15.21% (11.65%, 19.34%) 2.41% (-0.38%, 5.05%)
Table 2. Experimental validation for 3 incrementality tests. Each cell contains the point estimate and its corresponding 95% confidence intervals.
We compare DDA generated attribution $\hat{a}^{DDA}$ with ”ground truth” $\hat{a}^{\exp}$ and compute their difference, i.e. $\hat{d}^{DDA}=\hat{a}^{DDA}-\hat{a}^{\exp}$.
Confidence intervals of $\hat{a}^{\exp}$, $\hat{a}^{DDA}$, and $\hat{d}^{DDA}$ were derived using bootstrap sampling. We notice that the gaps between $\hat{a}^{\exp}$ and $\hat{a}^{DDA}$ are not statistically significant. We concluded that our DDA attention based attribution align with the experimental result.
4. Application to LinkedIn Ads Measurement
When applying LiDDA to the LinkedIn Ads use case, we also supply information about the advertiser, via a company embedding for the advertiser. This is concatenated with embeddings representing the member and their company. Similar to our GTM marketing model, we also include a touchpoint sequence representation (encoding the sequence of touch types, e.g., ”Onsite lead generation sponsored status update video ad impression”), touchpoint embeddings derived from campaign text, and positional encodings (representing the time between each touchpoint and the conversion). While GTM marketing has access to touchpoints across multiple channels, the Ads platform currently only has access to LinkedIn touchpoints. When applying the model to lead conversions obtained through LinkedIn’s Lead Generation form, on average the attribution weights show a time decay effect: touches closer to the conversion tend to get more credit (Figure 5). Meanwhile, We also observe the weight assigned to a specific touchpoint can vary, influenced by factors like the objective type, ad format, interaction type (impression, click), ad creative, days before conversion. We surface these insights at both path‐level detail and aggregated campaign/creative reports (Details in Appendix B.5).
Figure 5. Attribution weights for LinkedIn lead conversions show a time decay effect in aggregate; ad touches closer to the conversion tend to get more credit.
5. Conclusion
In this paper, we introduce LiDDA, a unified transformer-based data-driven attribution approach we developed and deployed at LinkedIn. We demonstrated the innovations to DDA methodology that combines an attention mechanism, positional embeddings, touchpoint imputation, and MMM calibration. We also shared learnings from implementation and deployment that are broadly applicable to the Marketing and Ad tech fields.
Appendix A Additional Details on Methods
A.1. Three‐Step Member Impression Allocation
Let $d$ index days, $c$ index paid channels, $j\in J_{i,c,d}$ index click‐events for member $i$, and $k\in\mathcal{P}_{d}$ index no‐click paths on day $d$. We observe external aggregates $(I^{\rm ext}_{c,d},\,C^{\rm ext}_{c,d})$, owned‐channel totals $(I^{\rm own}_{\rm total},C^{\rm own}_{\rm total},I^{\rm own}_{\rm clickers},C^{\rm own}_{\rm clickers})$, and a nonnegative weight $u_{k}$ for each no‐click path.
A.1.1. Member‐Only Impression Budget
Compute the number of impressions on channel $c$ and day $d$ attributable to known members by scaling the external impressions by the fraction of clicks from members:
$C^{\rm mem}_{c,d}=\sum_{i}\lvert J_{i,c,d}\rvert,\qquad I^{\rm mem}_{c,d}=I^{\rm ext}_{c,d}\,\frac{C^{\rm mem}_{c,d}}{C^{\rm ext}_{c,d}}.$
Summing over channels gives the total daily member‐impression budget:
$I_{d}\;=\;\sum_{c}I^{\rm mem}_{c,d}.$
A.1.2. Impute and Pre‐Allocate to Clicks
First, derive a down‐weight factor $\alpha<1$ from owned‐channel behavior:
$R^{\rm own}_{\rm all}=\frac{I^{\rm own}_{\rm total}}{C^{\rm own}_{\rm total}},\quad R^{\rm own}_{\rm clickers}=\frac{I^{\rm own}_{\rm clickers}}{C^{\rm own}_{\rm clickers}},\quad\alpha=\frac{R^{\rm own}_{\rm clickers}}{R^{\rm own}_{\rm all}}.$
Apply $\alpha$ to the raw external impressions‐per‐click to obtain the clickers’ expected impressions per click on $(c,d)$:
$R^{\rm ext}_{\rm all}=\frac{I^{\rm ext}_{c,d}}{C^{\rm ext}_{c,d}},\qquad r_{c,d}=\alpha\,R^{\rm ext}_{\rm all}\;\approx\;\mathbb{E}\bigl{[}\#\mathrm{imps}\mid\mathrm{click},c,d\bigr{]}.$
Then allocate to each click $j\in J_{i,c,d}$ at least one impression by either
[…]
and compute the pre‐allocated sum and leftover:
$I^{\rm pre}_{d}=\sum_{j}m_{j},\qquad L_{d}=I_{d}-I^{\rm pre}_{d}.$
A.1.3. Probabilistic Allocation of Leftover
Assign the remaining $L_{d}$ impressions across no‐click paths $k\in\mathcal{P}_{d}$ proportionally to weights $u_{k}$. Define
$p_{k}=\frac{u_{k}}{\sum_{\ell\in\mathcal{P}_{d}}u_{\ell}},\qquad\lambda_{k}=L_{d}\,p_{k}.$
Draw $\tilde{I}_{k}\sim\mathrm{Poisson}(\lambda_{k})$ independently, then rescale to exactly exhaust $L_{d}$:
$I_{k}=\Bigl{\lfloor}\tilde{I}_{k}\times\frac{L_{d}}{\sum_{\ell}\tilde{I}_{\ell}}\Bigr{\rfloor},\qquad\sum_{k}I_{k}=L_{d}.$
A.1.4. Result
Each click‐event $j$ receives $m_{j}\geq 1$ (on average $r_{c,d}$), each no‐click path $k$ receives $I_{k}\geq 0$, and
$\sum_{j}m_{j}\;+\;\sum_{k}I_{k}\;=\;I_{d},$
so the member‐impression budget is exactly allocated.
A.1.5. Remarks
We note that:
  • Both Geometric distribution and Poisson distribution can be appoximated by Normal distribution if needed;
  • $u_{k}\propto\mathbb{E}\bigl{[}\#\mathrm{imps}\mid\mathrm{noClick},k\bigr{]}$ can be assigned by contextual information such as corresponding campaign audience and location;
  • $r_{c,d}$ can be estimated with more complex methods, e.g., $r_{i,c,d}=f(x_{i},x_{c},x_{d})$ via a NN that leverage both member level data from owned channel and aggregate data from external channels.
A.1.6. Joint Convex‐Optimization Formulation
Instead of treating click allocation (Step 2) and leftover allocation (Step 3) separately, we can solve them in one convex‐optimization step.
Introduce: (i) Variables $x_{j}$ for the impressions assigned to click‐event $j\in J_{*,*,d}$; (ii) Variables $y_{k}$ for the impressions assigned to no‐click path $k\in\mathcal{P}_{d}$; (iii) Total budget $I_{d}$ from Step 1; (iv) Positive weights
[…]
reflecting the expected imps‐per‐click and the no‐click path priorities.
We then solve the following optimization:
$\underset{\{x_{j}\},\{y_{k}\}}{\text{maximize}}$ $\sum_{j\in J_{*,*,d}}w_{j}\,\log(x_{j})\;+\;\sum_{k\in\mathcal{P}_{d}}u_{k}\,\log(y_{k})$
subject to $\sum_{j\in J_{*,*,d}}x_{j}\;+\;\sum_{k\in\mathcal{P}_{d}}y_{k}\;=\;I_{d},$
[…]
KKT‐based solution. The Karush–Kuhn–Tucker conditions imply that at optimum there is a multiplier $\lambda>0$ such that
$x_{j}=\max\!\Bigl{(}1,\;\frac{w_{j}}{\lambda}\Bigr{)},\qquad y_{k}=\frac{u_{k}}{\lambda},$
and $\lambda$ must satisfy:
(8) $\sum_{j}\max\!\Bigl{(}1,\;\frac{w_{j}}{\lambda}\Bigr{)}+\sum_{k}\frac{u_{k}}{\lambda}=I_{d}.$
Eq. 8 can be solved by Algorithm 3, with complexity $O(J\log(J))$.
To summarize, the solution:
1. Honors the total budget: $\sum_{j}x_{j}+\sum_{k}y_{k}=I_{d}$.
2. Guarantees each click‐bucket $x_{j}\geq 1$ whenever $w_{j}>0$.
3. Allocates impressions in proportion to the expected click demand $w_{j}$ and the no‐click weights $u_{k}$.
By choosing $w_{j}=r_{c,d}$ and $u_{k}$ as in Steps 2–3, we recover a deterministic analogue of the previous stochastic Geometric+Poisson draws.
Algorithm 3 Compute Optimal $\lambda$ and Allocations
Input: Click‐weights $\{w_{j}\}_{j=1}^{J}$, path‐weights $\{u_{k}\}_{k=1}^{K}$, total budget $I_{d}$
Output: Impressions $\{x_{j}\},\{y_{k}\}$
Sort $w_{(1)}\geq w_{(2)}\geq\dots\geq w_{(J)}$
Compute $S_{u}\leftarrow\sum_{k=1}^{K}u_{k}$
$S_{w}\leftarrow\sum_{j=1}^{J}w_{(j)}$
for $m=0$ to $J$ do
     $S_{w}^{(m)}\leftarrow S_{w}-\sum_{j=1}^{m}w_{(j)}$
     $\lambda_{m}\leftarrow\frac{S_{w}^{(m)}+S_{u}}{\,I_{d}-m\,}$
end for
Find the unique $m^{*}$ such that
$w_{(m^{*})}\;\leq\;\lambda_{m^{*}}\quad\text{and}\quad w_{(m^{*}+1)}\;>\;\lambda_{m^{*}},$
with conventions $w_{(0)}=\infty$, $w_{(J+1)}=0$.
$\lambda\leftarrow\lambda_{m^{*}}$
for each click $j=1,\dots,J$ do
     $x_{j}\leftarrow\max\bigl{(}1,\;w_{j}/\lambda\bigr{)}$
end for
for each path $k=1,\dots,K$ do
     $y_{k}\leftarrow u_{k}/\lambda$
end for
[…]
@@ A.1.6. Joint Convex‐Optimization Formulation
A.2. Path Pre-processing
A.2.1. Sessionization
We group impressions from certain channels in a path as a single touchpoint within a session if they are from the same campaign on the same day, as detailed in Algorithm 4. This helps reduce noise, particularly from on-platform channels where a member may observe the same campaign repeatedly due to scrolling behavior. With this approach, we model a single event however we retain the downsampling links for post-model processing to reallocate credit equally across the grouped events.
A.2.2. Downsampling
We further address signal bias by adjusting the retention rate of specific channel, action combinations according to tunable downsampling rates. We obtain these rates using an independent analysis to estimate the comparable impact of each channel using signals such as spend and outcomes. With these ratios, we randomly ”drop” some touchpoints to yield the expected ratio as described in Algorithm 5. We observe that this pre-processing step yields more reliable attribution results and path compositions better aligned with aggregate data.